# Late-time X-ray Observations of the Jetted Tidal Disruption Event AT2022cmc: The Relativistic Jet Shuts Off

### v1

@@ Abstract
facilities: Chandra X-ray Observatory software: astropy (Astropy Collaboration et al., 2022), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), SAOImageDS9 (Joye & Mandel, 2003), scipy (Virtanen et al., 2020), CIAO (v4.13) (Fruscione et al., 2006),
@@ 1 Introduction
The discovery of two $\gamma$-ray transients by the Swift/Burst Alert Telescope (BAT) in 2011 provided the first unambiguous cases of relativistic jets in TDEs (Levan et al., 2011; Burrows et al., 2011; Bloom et al., 2011; Zauderer et al., 2011; Cenko et al., 2012). Following their discoveries, Swift J164449.3+573451 (hereafter Sw J1644+57) and Swift J2058.4+0516 (hereafter Sw J2058+05) were localized to the nuclei of distant galaxies at $z=0.354$ and $z=1.1853$ respectively, and were notably similar in their early time behavior in the radio, infrared, optical, and X-ray bands. In both cases, the early X-ray emission exhibited rapid variability on timescales as short as $\lesssim 500$ seconds (Brown et al., 2015; Mangano et al., 2016), and X-ray luminosities several orders of magnitude above the Eddington limit for a $\sim 10^{6}\ \rm M_{\odot}$ black hole, below the maximum allowed mass of $\sim 10^{8}\ \rm M_{\odot}$ for TDE flares. For Sw J1644+57, the discovery of bright radio-to-millimeter synchrotron emission independently established the presence of a collimated relativistic jet with a Lorentz factor of $\Gamma\sim$ few (Zauderer et al., 2011; Berger et al., 2012). Sw J1644+57’s X-ray luminosity subsequently declined following roughly a $t^{-5/3}$ power law decay (Mangano et al., 2016), as expected for fallback accretion in a TDE (Rees, 1988), while Sw J2058+05 had a steeper decay with $t^{-2.2}$. A third jetted TDE, Swift J1112.2--8238 (hereafter Sw J1112--82; $z=0.89$), was discovered in June 2011 (Brown et al., 2015), although limited data are available for this event.
@@
Over a decade since the discoveries of the first jetted TDEs, observations have revealed that relativistic jets in TDEs are extremely rare (Alexander et al., 2020). Compared to the rate of non-jetted TDEs ($\sim 10^{3}\ \rm Gpc^{-3}\ yr^{-1}$; Stone et al. 2020), TDEs that appear to power on-axis relativistic jets comprise less than $1\%$ of the TDE population (Sun et al., 2015; Andreoni et al., 2022a), suggesting that such events may require special conditions (e.g., high black hole spins or a strong magnetic flux threading the black hole; Tchekhovskoy et al. 2014) or implicate viewing angle effects in which the majority of jets are beamed out of our line of sight. The apparent diversity in the X-ray spectra of jetted and non-jetted TDEs (Auchettl et al., 2018), with the former exhibiting generally harder X-ray emission, may be naturally reconciled by the observer’s viewing angle with respect to the accretion disk (Dai et al., 2018). Alternatively, a misalignment between the black hole spin axis and orbital plane of the star may produce a quasi-spherical outflow that prevents all but the most powerful and highly magnetized jets from escaping, leading to the observed dichotomy (Lu et al., 2023; Teboul & Metzger, 2023).
@@
The discovery of the relativistic TDE AT2022cmc on 2022 February 11 by the Zwicky Transient Facility (ZTF) marked the first jetted TDE discovered in the optical and the first observed to launch a relativistic jet in the last decade. Unlike the Swift-detected events, AT2022cmc was initially discovered as a fast-fading optical transient (Andreoni et al., 2022b). Subsequent spectroscopic observations identified a redshift of $z=1.193$ (Tanvir et al., 2022), heralding AT2022cmc as the furthest TDE discovered to date. The early time X-ray emission from AT2022cmc was highly variable (Pasham et al., 2023), similar to the early X-ray light curve of Sw J1644+57, and was interpreted as synchrotron radiation due to energy dissipation within a magnetically dominated jet (Yao et al., 2024). Observations at radio and submillimeter wavelengths further supported the presence of a relativistic jet expanding into an ambient medium (Andreoni et al., 2022a; Pasham et al., 2023), as with Sw J1644+57.
@@
As one of a small sample of jetted TDEs, AT2022cmc affords a rare opportunity to gain insight into the emission mechanism operating in the accretion disk at early times (Pasham et al., 2023; Yao et al., 2024), as well as the properties of the SMBH and disrupted star through long-term monitoring of the X-ray light curve. Here we present late-time Chandra X-ray observations of AT2022cmc. Our observations demonstrate that the relativistic jet has shut off at $t_{\rm rest}\sim 100$ days, as evidenced by a sudden drop in the X-ray luminosity and a deviation from an earlier power-law decline. The X-ray data allow us to uniquely probe the mass of the disrupting SMBH and the total accreted mass. We adopt the discovery epoch MJD 59621.4463 (2022 February 11 at 10:42:40 UTC) as the reference epoch. Throughout the paper, we use the Planck cosmological parameters for a flat $\Lambda$CDM universe, with $H_{0}$ = 67.66 km s-1 Mpc-1, $\Omega_{m}=0.310$, and $\Omega_{\lambda}=0.690$ (Planck Collaboration et al., 2020).
@@ 3.1 The Relativistic Jet Shuts Off
Based on the results of our MCMC analysis, we find evidence for a break in the X-ray light curve at $t_{\rm off}\equiv t_{\rm rest}=98^{+38}_{-26}$ days ($t_{\rm obs}\approx 215$ days). Compared to Sw J2058+05 and Sw 1644+57, the two known jetted TDEs with observed X-ray drop-offs at $t_{\rm rest}\approx 200$ and $370$ days, respectively, the drop in AT2022cmc’s X-ray light curve occurs a factor of two to four times earlier. We explore the implications of this in Section 3.3.
@@ 3.2 Estimating the Mass of the Disrupting SMBH: X-ray Luminosity
Our results suggest that jetted TDEs may preferentially form from lower mass black holes when compared to non-relativistic events which comprise the bulk of the TDE population and typically occur in galaxies hosting more massive black holes (Mockler et al., 2019; Ryu et al., 2020; Nicholl et al., 2022; Hammerstein et al., 2023). If the latter form in systems with larger black hole masses, the paucity of jets may be a direct consequence of the lower jet efficiency and higher disk efficiency in this mass regime. Indeed, a positive correlation between black hole mass and disk radiative efficiency has been suggested in a sample of optically bright TDEs (Nicholl et al., 2022). At high black hole masses, such systems would therefore require extremely powerful jets (i.e., large jet efficiences) to outshine their accretion disks.
@@ 3.3 Estimating the Mass of the Disrupting SMBH: Timescale of Jet Shut-off
Using equations 6, 9, 10, and 11 with $\alpha=2.1$ as constrained from our MCMC analysis, our inferred jet shut-off time of $t_{\rm rest}=98$ days, and assuming a solar mass star, in Figure 3 we plot black hole mass as a function of $f_{\rm in}$, where we adopt $\epsilon_{\rm disk}=0.1$ as a typical radiative efficiency. [^6] In reality, the radiative efficiency for super-Eddington accretion disks decreases as the Eddington ratio $\lambda$ increases (Jiang et al., 2019). For our purposes, and given that we carry out our calculations at the time of jet shut-off when the radiative efficiency is expected to be close to the typical value, we adopt a constant radiative efficiency $\epsilon_{\rm disk}=0.1$. The result is a curve in the $f_{\rm in}-M_{\rm BH}$ phase-space corresponding to the fraction of accreted mass relative to a full disruption, where $f_{\rm in}=1$ implies a complete disruption in which 0.5 $M_{\odot}$ is accreted onto the black hole. [^7] Here we assume the tidal disruption of a $1\ M_{\odot}$ star in which half of the stellar material is unbound from the system and additionally allow for the possibility that the star is only partially disrupted.
@@
Given $L_{\rm jet,off}$ and $t_{\rm off}$, for fixed values of $M_{\rm BH}$ in Figure 3, there exist corresponding values of $f_{\rm in}$ and $\eta_{\rm jet}$. Under the above supposition that the jet and disk luminosities are comparable (i.e., $\eta_{\rm jet}\gtrsim 10^{-1}$), as in powerful blazars and radio-loud quasars, and $M_{\rm BH}\lesssim 10^{5}\ \rm M_{\odot}$, $f_{\rm in}$ is constrained to $\lesssim 10^{-2}$ for AT2022cmc, implying a small fraction of total accreted mass relative to a complete disruption in which $0.5\ \rm M_{\odot}$ is accreted. This is consistent with the framework presented by Metzger & Stone 2016 in which they postulate that $f_{\rm in}\ll 1$ is required to produce the observed outflows in TDEs. On the other hand, if the jet-disk coupling in TDEs is intrinsically distinct from that of AGN, the jet may be radiatively inefficient ($\eta_{\rm jet}\lesssim 10^{-3}$) at high black hole masses ($M_{\rm BH}\sim 10^{7}\ \rm M_{\odot}$) despite a large fraction of accreted mass ($f_{\rm in}\gtrsim 0.1$). In other words, the X-ray luminosity would represent a small fraction of the total energy budget of the system. Such a scenario may arise due to differences in the available magnetic reservoir for TDEs (Kelley et al., 2014), for example, or if a large fraction of jets are choked (Lu et al., 2023; Teboul & Metzger, 2023) or launched off-axis (Dai et al., 2018). Indeed, the latter scenario may be reflected by the smaller fraction of relativistic jetted TDEs relative to the fraction of radio-quiet quasars (Alexander et al., 2020).
@@
Adopting the same methodology for Sw J1644+57888We adopt $\alpha=5/3$ for Sw J1644+57, consistent with its observed X-ray light curve (Burrows et al., 2011). and Sw J2058+05999We constrain the timescale of jet shut-off and the power-law index for Sw J2058+05 in Appendix A., we trace out similiar regions in the $f_{\rm in}-M_{\rm BH}$ phase-space in Figure 3. We note that in both cases, the timescale of jet shut-off occurs a factor of two to four times later than in AT2022cmc; requiring the same condition of $\eta_{\rm jet}\gtrsim 10^{-1}$ therefore allows for somewhat larger black hole masses relative to AT2022cmc of $\sim$ a few ×$\times 10^{5}\ \rm M_{\odot}$ and hence a larger fraction of accreted mass $f_{\rm in}\lesssim 0.1$, implying that the total accreted mass in these systems is more comparable to that of a complete disruption.
@@ 3.4 Comparison to Other SMBH Mass Estimates
We compare our SMBH mass estimates for AT2022cmc and other jetted TDEs as derived from the jet shut-off time and X-ray luminosity to estimates derived via other methods. First, we use the observed variability in the X-ray light curves at early times to place an upper limit on the mass of the SMBH. In particular, we can equate the size of the X-ray emitting region $r_{s}$ to the Schwarzschild radius of a black hole with mass $M_{\rm BH}$. For an observed variability timescale $t_{\rm var,sec}\sim r_{s}/c$ in seconds, [^10] We note that beaming should increase the black hole mass by a factor $\approx\Gamma_{j}^{2}$, where $\Gamma_{j}$ is the Lorentz factor of the jet. Here we neglect beaming in light of large uncertainties on $\Gamma_{j}$. the mass of the SMBH is given by
@@
For Sw J1644+57, the early time X-ray light curve out to $t_{\rm obs}\approx 10$ days is punctuated by rapid variability on timescales as short as $\sim 100$ s (Mangano et al., 2016). We therefore constrain the black hole mass for Sw J1644+57 based on variability arguments to $M_{BH}\lesssim 8\times 10^{6}\ \rm M_{\odot}$, as in Burrows et al. (2011). Similarly, the X-ray light curve of Sw J1112--82 exhibited variability on timescales of a few thousand seconds (Brown et al., 2015), corresponding to a limit on the SMBH mass of $\lesssim 2\times 10^{7}\ \rm M_{\odot}$. Finally, for Sw J2058+05, the observed variability timescale of $\sim 500$ s constrains the SMBH mass to $\lesssim 2\times 10^{8}\ \rm M_{\odot}$. This is above the maximum allowed mass for TDEs of $<10^{8}\ \rm M_{\odot}$, above which the disruption radius lies within the Schwarzschild radius and the star is thus swallowed whole (Rees, 1988), and hence is not constraining.

### v2

@@ Abstract
facilities: Chandra X-ray Observatory software: Astropy (Astropy Collaboration et al., 2013, 2018, 2022), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), SAOImageDS9 (Joye & Mandel, 2003), scipy (Virtanen et al., 2020), CIAO (v4.13) (Fruscione et al., 2006),
@@ 1 Introduction
The discovery of two $\gamma$-ray transients by the Swift/Burst Alert Telescope (BAT) in 2011 provided the first unambiguous cases of relativistic jets in TDEs (Levan et al., 2011; Burrows et al., 2011; Bloom et al., 2011; Zauderer et al., 2011; Cenko et al., 2012). Following their discoveries, Swift J164449.3+573451 (hereafter Sw J1644+57) and Swift J2058.4+0516 (hereafter Sw J2058+05) were localized to the nuclei of distant galaxies at $z=0.354$ and $z=1.1853$ respectively, and were notably similar in their early time behavior in the radio, infrared, and X-ray bands. In both cases, the early X-ray emission exhibited rapid variability on timescales as short as $\lesssim 500$ seconds (Brown et al., 2015; Mangano et al., 2016), and X-ray luminosities several orders of magnitude above the Eddington limit for a $\sim 10^{6}\ \rm M_{\odot}$ black hole, below the maximum allowed mass of $\sim 10^{8}\ \rm M_{\odot}$ for TDE flares. For Sw J1644+57, the discovery of bright radio-to-millimeter synchrotron emission independently established the presence of a collimated relativistic jet with a Lorentz factor of $\Gamma\sim$ few (Zauderer et al., 2011; Berger et al., 2012). Sw J1644+57’s X-ray luminosity subsequently declined following roughly a $t^{-5/3}$ power law decay (Mangano et al., 2016), as expected for fallback accretion in a TDE (Rees, 1988), while Sw J2058+05 had a steeper decay with $t^{-2.2}$. A third jetted TDE, Swift J1112.2--8238 (hereafter Sw J1112--82; $z=0.89$), was discovered in June 2011 (Brown et al., 2015), although limited data are available for this event.
@@
Over a decade since the discoveries of the first jetted TDEs, observations have revealed that relativistic jets in TDEs are extremely rare (Alexander et al., 2020). Compared to the rate of non-jetted TDEs ($\sim 10^{3}\ \rm Gpc^{-3}\ yr^{-1}$; Stone et al. 2020), TDEs that appear to power on-axis relativistic jets comprise less than $1\%$ of the TDE population (Sun et al., 2015; Andreoni et al., 2022a), suggesting that such events may require special conditions (e.g., high black hole spins or a strong magnetic flux threading the black hole; Tchekhovskoy et al. 2014) or implicate viewing angle effects in which the majority of jets are beamed out of our line of sight. Indeed, the emergence of delayed radio emission from a subset of TDEs (e.g., Horesh et al. 2021a; Cendes et al. 2023) can potentially be explained by off-axis jets (Sfaradi et al., 2024; Sato et al., 2024), though non-relativistic outflows provide a viable mechanism as well. Nonetheless, the fraction of powerful jets similar to Sw J1644+57 still appears to be extremely small. The apparent diversity in the X-ray spectra of jetted and non-jetted TDEs (Auchettl et al., 2018), with the former exhibiting generally harder X-ray emission, may be naturally reconciled by the observer’s viewing angle with respect to the accretion disk (Dai et al., 2018). Alternatively, a misalignment between the black hole spin axis and orbital plane of the star may produce a quasi-spherical outflow that prevents all but the most powerful and highly magnetized jets from escaping, leading to the observed dichotomy (Lu et al., 2023; Teboul & Metzger, 2023).
@@
The discovery of the relativistic TDE AT2022cmc on 2022 February 11 by the Zwicky Transient Facility (ZTF) marked the first jetted TDE discovered in the optical and the first observed to launch a relativistic jet in the last decade. Unlike the Swift-detected events, AT2022cmc was initially discovered as a fast-fading optical transient (Andreoni et al., 2022b). Subsequent spectroscopic observations identified a redshift of $z=1.193$ (Tanvir et al., 2022), heralding AT2022cmc as the furthest TDE discovered to date. The early time X-ray emission from AT2022cmc was highly variable (Pasham et al., 2023), similar to the early X-ray light curve of Sw J1644+57, and was interpreted as synchrotron radiation due to energy dissipation within a magnetically dominated jet (Yao et al., 2024). Observations at radio and submillimeter wavelengths further supported the presence of a relativistic jet expanding into an ambient medium (Andreoni et al., 2022a; Pasham et al., 2023; Rhodes et al., 2023), as with Sw J1644+57.
@@
As one of a small sample of jetted TDEs, AT2022cmc affords a rare opportunity to gain insight into the emission mechanism operating in the accretion disk at early times (Pasham et al., 2023; Yao et al., 2024), as well as the properties of the SMBH and disrupted star through long-term monitoring of the X-ray light curve. Here we present late-time Chandra X-ray observations of AT2022cmc. Proceeding under the assumption that the X-ray emission from AT2022cmc is dominated by a relativistic jet, our observations demonstrate that the jet has shut off at $t_{\rm rest}\sim 100$ days, as evidenced by a sudden drop in the X-ray luminosity and a deviation from an earlier power-law decline. The X-ray data allow us to uniquely probe the mass of the disrupting SMBH and the total accreted mass. We adopt the discovery epoch MJD 59621.4463 (2022 February 11 at 10:42:40 UTC) as the reference epoch. Throughout the paper, we use the Planck cosmological parameters for a flat $\Lambda$CDM universe, with $H_{0}$ = 67.66 km s-1 Mpc-1, $\Omega_{m}=0.310$, and $\Omega_{\lambda}=0.690$ (Planck Collaboration et al., 2020).
@@ 3.1 The Relativistic Jet Shuts Off
Based on the results of our MCMC analysis, we find evidence for a break in the X-ray light curve at $t_{\rm off}\equiv t_{\rm rest}=98^{+38}_{-26}$ days ($t_{\rm obs}\approx 215$ days). We note a degeneracy between the jet shut-off time $t_{\rm off}$ and the flux normalization $F_{X}$ given the sparse temporal coverage in the light curve between the last detection and our late-time Chandra limits. Compared to Sw J2058+05 and Sw J1644+57, the two known jetted TDEs with observed X-ray drop-offs at $t_{\rm rest}\approx 200$ and $370$ days, respectively, the drop in AT2022cmc’s X-ray light curve occurs a factor of two to four times earlier. We explore the implications of this in Section 3.3.
@@ 3.2 Estimating the Mass of the Disrupting SMBH: X-ray Luminosity
Our results suggest that jetted TDEs may preferentially form from lower mass black holes when compared to non-relativistic events which comprise the bulk of the TDE population and typically occur in galaxies hosting more massive black holes (Mockler et al., 2019; Ryu et al., 2020; Nicholl et al., 2022; Hammerstein et al., 2023), similar to the findings of Somalwar et al. 2023 based on a sample of radio-selected TDEs. If the latter form in systems with larger black hole masses, the paucity of jets may be a direct consequence of the lower jet efficiency and higher disk efficiency in this mass regime. Indeed, a positive correlation between black hole mass and disk radiative efficiency has been suggested in a sample of optically bright TDEs (Nicholl et al., 2022). At high black hole masses, such systems would therefore require extremely powerful jets (i.e., large jet efficiencies) to outshine their accretion disks.
@@ 3.3 Estimating the Mass of the Disrupting SMBH: Timescale of Jet Shut-off
Using equations 6, 9, 10, and 11 with $\alpha=2.1$ as constrained from our MCMC analysis, our inferred jet shut-off time of $t_{\rm rest}=98$ days, and assuming a solar mass star, in Figure 3 we plot black hole mass as a function of $f_{\rm in}$, where we adopt $\epsilon_{\rm disk}=0.1$ as a typical radiative efficiency. [^7] In reality, the radiative efficiency for super-Eddington accretion disks decreases as the Eddington ratio $\lambda$ increases (Jiang et al., 2019). For our purposes, and given that we carry out our calculations at the time of jet shut-off when the radiative efficiency is expected to be close to the typical value, we adopt a constant radiative efficiency $\epsilon_{\rm disk}=0.1$. The result is a curve in the $f_{\rm in}-M_{\rm BH}$ phase-space corresponding to the fraction of accreted mass relative to a full disruption, where $f_{\rm in}=1$ implies a complete disruption in which 0.5 $M_{\odot}$ is accreted onto the black hole. [^8] Here we assume the tidal disruption of a $1\ M_{\odot}$ main sequence star in which half of the stellar material is unbound from the system as the most plausible scenario and additionally allow for the possibility that the star is only partially disrupted. In the case of a white dwarf disruption, which is tenable for a $10^{5}\ \rm M_{\odot}$ black hole and indeed was suggested for Sw J1644+57 based on the short timescale variability in the Swift/XRT lightcurve (Krolik & Piran, 2011), in which the radius and mass of the star are inversely related ($r_{*}\propto M_{*}^{-1/3}$), the mass accretion rate on the same timescale is approximately $10^{3}$ times smaller (Tchekhovskoy et al., 2014), implying a larger accreted mass fraction for a fixed black hole mass.
@@
Given $L_{\rm jet,off}$ and $t_{\rm off}$, for fixed values of $M_{\rm BH}$ in Figure 3, there exist corresponding values of $f_{\rm in}$ and $\eta_{\rm jet}$. Under the above supposition that the jet and disk luminosities are comparable (i.e. $\eta_{\rm jet}\gtrsim 10^{-1}$), as in powerful blazars and radio-loud quasars, and $M_{\rm BH}\lesssim 10^{5}\ \rm M_{\odot}$, $f_{\rm in}$ is constrained to $\lesssim 10^{-2}$ for AT2022cmc, implying a small fraction of total accreted mass relative to a complete disruption in which $0.5\ \rm M_{\odot}$ is accreted. This is consistent with the framework presented by Metzger & Stone 2016 in which they postulate that $f_{\rm in}\ll 1$ is required to produce the observed outflows in TDEs. On the other hand, if the jet-disk coupling in TDEs is intrinsically distinct from that of AGN, the jet may be radiatively inefficient ($\eta_{\rm jet}\lesssim 10^{-3}$) at high black hole masses ($M_{\rm BH}\sim 10^{7}\ \rm M_{\odot}$) despite a large fraction of accreted mass ($f_{\rm in}\gtrsim 0.1$). In other words, the X-ray luminosity would represent a small fraction of the total energy budget of the system. Such a scenario may arise due to differences in the available magnetic reservoir for TDEs (Kelley et al., 2014), for example, or if a large fraction of jets are choked (Lu et al., 2023; Teboul & Metzger, 2023) or launched off-axis (Dai et al., 2018). Indeed, the latter scenario may be reflected by the smaller fraction of relativistic jetted TDEs relative to the fraction of radio-quiet quasars (Alexander et al., 2020).
@@
Adopting the same methodology for Sw J1644+57999We adopt $\alpha=5/3$ for Sw J1644+57, consistent with its observed X-ray light curve (Burrows et al., 2011). and Sw J2058+05101010We constrain the timescale of jet shut-off and the power-law index for Sw J2058+05 in Appendix A., we trace out similiar regions in the $f_{\rm in}-M_{\rm BH}$ phase-space in Figure 3. We note that in both cases, the timescale of jet shut-off occurs a factor of two to four times later than in AT2022cmc; requiring the same condition of $\eta_{\rm jet}\gtrsim 10^{-1}$ therefore allows for somewhat larger black hole masses relative to AT2022cmc of $\sim$ a few ×$\times 10^{5}\ \rm M_{\odot}$ and hence a larger fraction of accreted mass $f_{\rm in}\lesssim 0.1$, implying that the total accreted mass in these systems is more comparable to that of a complete disruption.
@@ 3.4 Comparison to Other SMBH Mass Estimates
We compare our SMBH mass estimates for AT2022cmc and other jetted TDEs as derived from the jet shut-off time and X-ray luminosity to estimates derived via other methods. First, we use the observed variability in the X-ray light curves at early times to place an upper limit on the mass of the SMBH. In particular, we can equate the size of the X-ray emitting region $r_{s}$ to the Schwarzschild radius of a black hole with mass $M_{\rm BH}$. For an observed variability timescale $t_{\rm var,sec}\sim r_{s}/c$ in seconds, the mass of the SMBH is given by
@@
For Sw J1644+57, the early time X-ray light curve out to $t_{\rm obs}\approx 10$ days is punctuated by rapid variability on timescales as short as $\sim 100$ s (Mangano et al., 2016). We therefore constrain the black hole mass for Sw J1644+57 based on variability arguments to $M_{BH}\lesssim 8\times 10^{6}\ \rm M_{\odot}$, as in Burrows et al. (2011). Conversely, the putative detection of quasi-periodic oscillations (QPOs) in the X-ray light curve of Sw J1644+57 (Reis et al., 2012) enabled constraints on the black hole mass of $10^{5}-10^{6}\ \rm M_{\odot}$ (Abramowicz & Liu, 2012), comparable to the black holes mass we infer in our jet shut-off paradigm. For Sw J1112--82, the X-ray light curve exhibited variability on timescales of a few thousand seconds (Brown et al., 2015), corresponding to a limit on the SMBH mass of $\lesssim 2\times 10^{7}\ \rm M_{\odot}$. Finally, for Sw J2058+05, the observed variability timescale of $\sim 500$ s constrains the SMBH mass to $\lesssim 2\times 10^{8}\ \rm M_{\odot}$. This is above the maximum allowed mass for TDEs of $<10^{8}\ \rm M_{\odot}$, above which the disruption radius lies within the Schwarzschild radius and the star is thus swallowed whole (Rees, 1988), and hence is not constraining.